what is the area of a square whose perimeter is 76 dm?
Found 2 solutions by cleomenius, amoresroy:
Answer by cleomenius(959)

The first step is to obtain the length of a side.
P = s + s + s + s
78 = 19 + 19 + 19 + 19
The formula for the area of a square is s^2.
Therefore, your area will be 19^2 = 361
Cleomenius.

Answer by amoresroy(361)

what is the area of a square whose perimeter is 76 dm?
Let x = the side of the square
4x = the perimeter of the square
4x = 76
x = 19
x^2 = area of square
= 19*19
= 361 dm^2
The area of square is 361 square dm.